# Adapting coherent-state superpositions in noisy channels

### v1

@@ 1 Introduction
The main challenge faced by the optical systems is the decoherence, mainly optical loss and thermal noise. While high quality quantum states may offer the possibility of error correction gottesman2001,menicucci2014,baragiola2019,pantaleoni2020, decoherence poses severe limits on the preparation and propagation of such states yukawa2013,eaton2022. While energy loss can never completely remove certain quantum non-Gaussian features from quantum states straka2014, losing half of the signal to the environment is enough to completely obliterate any negativity of the Wigner function, thus preventing any possibility of quantum advantage mari2012. Loss combined with noise present in the environment makes its decay even faster.
It has been established that the effects of loss in photon qubit systems can be probabilistically mitigated by suitable pre-processing of optical signals micuda2012,gagatsos2014,chrzanowski2014 or channel engineering starek2020. Most auspiciously, CS states specifically can be made more resilient to loss by applying a suitable deterministic squeezing operation filip2001,serafini2004,teh2020,filip2013,lejeannic2018,brewster2018,pan2023. Specifically, it has been shown for pure lossy channels in lejeannic2018. However, to what extent can these methods be used for mitigation of added channel noise together with loss has not been studied so far.
To address this question we consider a CS state propagating through a general series of channels incorporating loss and added phase-sensitive noise. We derive a general condition that needs to be satisfied to preserve any negativity present in the Wigner function of the quantum state based on optimal deterministic squeezing operations, that maximizes the central negativity of the Wigner function. In addition we extend the general condition to include even-parity CS states in the form of a necessary condition that must be satisfied for any negative values of the Wigner function to exist.
Our analysis based on negativity of the Wigner function is complemented by adapting the squeezing operations to maximize the Hilbert-Schmidt distance between CS states of opposite parities transmitted through the decohering channels.
2 Minimizing decoherence in a single noisy channel
@@
where we set ${\sqrt{2}\xi\equiv X_{0}+\imath P_{0}}$ and therefore ${2\lvert\xi\rvert^{2}\equiv X_{0}^{2}+P_{0}^{2}}$. As for the final building block, omitted in (3), we have ${w_{-\xi,\xi}(X,P)\equiv\operatorname{conj}\{w_{\xi,-\xi}(X,P)\}}$. These two complex-exponential blocks add up to a real-valued cosine.
@@
$f_{X}$ $=\sqrt{\eta}\mathrm{e}^{-\gamma},$ $\sigma_{X}$ $=2(1-\eta)\mathrm{e}^{-2\gamma_{t}}V\;\text{,}$ (8)
$f_{P}$ $=\sqrt{\eta}\mathrm{e}^{+\gamma},$ $\sigma_{P}$ $=2(1-\eta)\mathrm{e}^{+2\gamma_{t}}V\;\text{.}$
@@
$W_{\xi,-}^{\prime}(0,0)=f_{\xi,-}\Bigg{\{}\exp\left(-\frac{f_{X}^{2}}{V_{X}}X_{0}^{2}-\frac{f_{P}^{2}}{V_{P}}P_{0}^{2}\right)-\exp\left(-\frac{\sigma_{X}}{V_{X}}P_{0}^{2}-\frac{\sigma_{P}}{V_{P}}X_{0}^{2}\right)\Bigg{\}}$ (9)
where the normalization factor reads ${f_{\xi,-}^{-1}=\pi\sqrt{V_{X}V_{P}}\left(1-\exp(-2\lvert\xi\rvert^{2})\right)}$. It will be negative iff
@@
This criterion can be generalized to include even-parity states at the cost of sufficiency. If the transformed Wigner function (7) of the transmitted CS state is to attain negative values anywhere, the channel must necessarily satisfy ${f_{X}^{2}f_{P}^{2}-\sigma_{X}\sigma_{P}>0}$. In particular,
$\exists(X,P)\in\mathbb{R}^{2}:W^{\prime}_{\xi,\pm}(X,P)<0\implies f_{X}^{2}f_{P}^{2}-\sigma_{X}\sigma_{P}>0\;\text{.}$ (11)
This condition is both necessary and sufficient for the central negativity of odd-parity CS states. Alternatively the necessary condition, equivalent to (11),
$f_{X}^{2}f_{P}^{2}-\sigma_{X}\sigma_{P}<0\implies\forall(X,P)\in\mathbb{R}^{2}:W_{\xi,\pm}(X,P)>0$ (12)
precludes negativity anywhere within the Wigner function (7). Both conditions are independent of the magnitude and parity of the transmitted CS state and depend only on the parameters of the channel. Both statements are proven and discussed within the supplementary material.
@@
3 Minimizing decoherence across multiple noisy channels
@@
$f_{X}$ $=\sqrt{\eta\eta^{\prime}}\mathrm{e}^{-(\gamma+\gamma^{\prime})}\;\text{,}$ (14)
$\sigma_{X}$ $=2(1-\eta^{\prime})\mathrm{e}^{-2\gamma^{\prime}_{t}}V^{\prime}+2\eta^{\prime}(1-\eta)\mathrm{e}^{-2(\gamma^{\prime}+\gamma_{t})}V\;\text{,}$
$f_{P}$ $=\sqrt{\eta\eta^{\prime}}\mathrm{e}^{+(\gamma+\gamma^{\prime})}\;\text{,}$
$\sigma_{P}$ $=2(1-\eta^{\prime})\mathrm{e}^{+2\gamma^{\prime}_{t}}V^{\prime}+2\eta^{\prime}(1-\eta)\mathrm{e}^{+2(\gamma^{\prime}+\gamma_{t})}V^{\prime}\;\text{,}$
@@
In Figure 5 we explore the central negativity of the odd-parity CS state with ${\sqrt{2}\xi=3}$ transmitted through a composite channel (14). Without loss of generality we set the transmittances in both channels identical to keep the visualisation two dimensional. The plots show the best attainable negativity and the optimal pre-squeezing and mid-squeezing rates. We consider a number of asymmetric thermal environments. In Figure 5 (a) the dashed lines represent the attainable CN without any protective squeezing operations applied. The solid black line represents the best attainable CN when optimal pre-squeezing and mid-squeezing optimal pre-squeezing and mid-squeezing operations are applied to the CS state. The optimal CN depends only on the symmetric variance of the noise, rather than its asymmetry. The solid black line can be equally obtained for the same channel where both environments are symmetric. The solid black line in Figure 5 (b) represents the optimal pre-squeezing rate $\gamma$ obtained when there is no asymmetry in either thermal environment, whereas the colored lines correspond to different asymmetries. We recognize that these rates are constantly shifted from the symmetric case, the offset equal to $\gamma_{t}$ of the first channel. The colored bullets, obtained by adding the respective $\gamma_{t}$ rates to the black line, are used to highlight this observation. Furthermore, in the third plot, Figure 5 (c), we observe that the optimal mid-squeezing rates $\gamma^{\prime}$ match the differences ${\gamma^{\prime}\equiv\gamma^{\prime}_{t}-\gamma_{t}}$ between the asymmetries of both environments.
@@
4 Hilbert-Schmidt distance as a measure of state adaptation
@@ Acknowledgements
The authors acknowledge early discussions with H. Le Jeannic. J.P. acknowledges fruitful discussions with O. Solodovnikova. J.P. acknowledges using the computational cluster at the Department of Optics at Palacký University and using several open-source software libraries harris2020,virtanen2020,dalcin2021,hunter2007,meurer2017 in the computation and subsequent evaluation of presented results.
@@
Jan Provazník1$\star$ , Petr Marek1 , Julien Laurat2 , Radim Filip1
1 Department of Optics, Palacký University, 17. listopadu 1192/12, 771 46 Olomouc, Czech Republic
2 Laboratoire Kastler Brossel, Sorbonne Université, CNRS, ENS-Université PSL, Collège de France, 4 place Jussieu, 75005 Paris, France
$\star$ provaznik@optics.upol.cz
S.1 Conditions for presence of negativity in Wigner functions
It convenient to rewrite the complex-exponential form of the output Wigner function of the CS state into its manifestly real form. The coefficients $f_{X},f_{P},\sigma_{X},\sigma_{P},V_{X}$ and $V_{P}$ are all real and positive.
@@
where the factors $f_{\pm}$ stand for the normalization, given as
$f_{\pm}^{-1}=2\pi\sqrt{V_{X}V_{P}}\left(1\pm\exp(-X_{0}^{2}-P_{0}^{2})\right)\;\text{.}$ (S.2)
Let us begin with the sufficient and necessary condition for the existence of any negative area in the Wigner function of the odd-parity CS state. It is only natural to consider the central negativity in this particular case. The formula (S.1) simplifies into
@@
that is, if and only if, the second exponential is greater than the first one. Comparing their arguments directly reveals the condition
$-X_{0}^{2}\left(\frac{f_{X}f_{P}-\sigma_{X}\sigma_{P}}{V_{X}V_{P}}\right)-P_{0}^{2}\left(\frac{f_{X}f_{P}-\sigma_{X}\sigma_{P}}{V_{X}V_{P}}\right)<0\iff f_{X}f_{P}-\sigma_{X}\sigma_{P}>0\;\text{.}$ (S.5)
The central point of the odd-parity CS state will be negative if and only if ${f_{X}f_{P}-\sigma_{X}\sigma_{P}>0}$. Notably this inequality does not depend on the state. It depends solely on the channel. This concludes the proof for the odd-parity state. $\blacksquare$
The condition for even-parity states a slightly harder to obtain. Once again, we note that we are only interested in a necessary condition for existence of negativity anywhere within the Wigner function. In other words, we we are looking for a condition that, if violated, will preclude existence of any negativity within the Wigner function. This allows us to be somewhat liberal in the manipulation of the expression (S.1).
@@
±$\pm\cos\left(2\left[\frac{f_{P}}{V_{P}}X_{0}P^{\prime}-\frac{f_{X}}{V_{X}}P_{0}X^{\prime}\right]\right)\equiv-\left\lvert\cos\left(2\left[\frac{f_{P}}{V_{P}}X_{0}P^{\prime}-\frac{f_{X}}{V_{X}}P_{0}X^{\prime}\right]\right)\right\rvert\;\text{.}$ (S.6)
The rest of the proof assumes the Wigner function (S.1) is evaluated only at points of the phase space that satisfy this condition. With this assumption in mind, we can rewrite the curly braced expression from (S.1) into
$\exp(A)\cosh(B)-\exp(C)\left\lvert\cos(D)\right\rvert<0$ (S.7)
where we have substituted the arguments of the exponential, geometric and hyperbolic functions for simplicity. Since the hyperbolic cosine is always positive and lower bound by $1$, we have
$\exp(A)\cosh(B)-\exp(C)\left\lvert\cos(D)\right\rvert<0$ $\iff$ (S.8)
$\exp(A)<\exp(C)\frac{\left\lvert\cos(D)\right\rvert}{\cosh(B)}$ $\iff$
$\exp(A)<\exp(C)$ $\iff$
$-X_{0}^{2}\left(\frac{f_{X}f_{P}-\sigma_{X}\sigma_{P}}{V_{X}V_{P}}\right)-P_{0}^{2}\left(\frac{f_{X}f_{P}-\sigma_{X}\sigma_{P}}{V_{X}V_{P}}\right)<0$ $\iff$
$\boxed{f_{X}f_{P}-\sigma_{X}\sigma_{P}>0}\;\text{.}$
Where we obtained the third inequality $\exp(A)<\exp(C)$ as an upper bound on the right hand side of the second one. The condition ${f_{X}f_{P}-\sigma_{X}\sigma_{P}>0}$ then follows. This condition, while valid for both even and odd parity CS states, is merely a necessary condition for the existence of some negative point. In other words, if violated, there will be no negative area at all. $\blacksquare$
S.2 Closed form formula for the Hilbert-Schmidt distance
@@
$O_{\xi}$ $=2\pi\iint\limits_{\mathbb{R}^{2}}W_{\xi,+}(X,P)W_{\xi,-}(X,P)\mathrm{d}{X}\mathrm{d}{P}$

### v2

@@ Introduction
The main challenge faced by the optical systems is the decoherence, mainly due to optical loss and thermal noise. While high quality quantum states may offer the possibility of error correction gottesman2001,menicucci2014,baragiola2019,pantaleoni2020, decoherence poses severe limits on the preparation and propagation of such states yukawa2013,eaton2022. While energy loss can never completely remove certain quantum non-Gaussian features from quantum states straka2014, losing half of the signal to the environment is enough to completely obliterate any negativity of the Wigner function, thus preventing any possibility of quantum advantage mari2012. Loss combined with noise present in the environment makes its decay even faster.
It has been established that the effects of loss in photonic qubit systems can be probabilistically alleviated by suitable pre-processing of optical signals micuda2012,gagatsos2014,chrzanowski2014 or channel engineering starek2020. In particular, CS states can be made more resilient to loss by applying a deterministic squeezing operation filip2001,serafini2004,teh2020,filip2013,lejeannic2018,brewster2018,pan2023. Specifically, it has been shown for pure lossy channels in lejeannic2018. However, to what extent can these methods be used for mitigation of added channel noise together with loss has not been studied so far.
To address this question we consider a CS state propagating through a general series of channels incorporating loss and added phase-sensitive noise. We derive a general condition that needs to be satisfied to preserve any negativity present in the Wigner function of the quantum state based on optimal deterministic squeezing operations, that maximizes the central negativity of the Wigner function. In addition we extend the general condition to include even-parity CS states. This condition that must be satisfied for any negative values of the Wigner function to exist.
Our analysis based primarily on negativity of the Wigner function is complemented by adapting the squeezing operations to maximize the Hilbert-Schmidt distance between CS states of opposite parities transmitted through the decohering channels.
Minimizing decoherence in a single noisy channel
@@
where we set ${\sqrt{2}\xi\equiv X_{0}+\imath P_{0}}$ and therefore ${2\lvert\xi\rvert^{2}\equiv X_{0}^{2}+P_{0}^{2}}$. As for the final building block, omitted in (3), we have ${w_{-\xi,\xi}(X,P)\equiv{\{}^{*}w_{\xi,-\xi}(X,P)\}}$. These two complex-exponential blocks add up to a real-valued cosine.
@@
$f_{X}$ $=\sqrt{\eta}\mathrm{e}[-\gamma],$ $\sigma_{X}$ $=2(1-\eta)\mathrm{e}[-2\gamma_{t}]V\;\text{,}$ (8)
$f_{P}$ $=\sqrt{\eta}\mathrm{e}[+\gamma],$ $\sigma_{P}$ $=2(1-\eta)\mathrm{e}[+2\gamma_{t}]V\;\text{.}$
@@
$W_{\xi,-}^{\prime}(0,0)=m_{\xi,-}\Bigg{\{}\exp\left(-\frac{f_{X}^{2}}{V_{X}}X_{0}^{2}-\frac{f_{P}^{2}}{V_{P}}P_{0}^{2}\right)-\exp\left(-\frac{\sigma_{X}}{V_{X}}P_{0}^{2}-\frac{\sigma_{P}}{V_{P}}X_{0}^{2}\right)\Bigg{\}}$ (9)
where the normalization factor reads ${m_{\xi,-}^{-1}=\pi\sqrt{V_{X}V_{P}}\left(1-\exp(-2\lvert\xi\rvert^{2})\right)}$. It will be negative iff
@@
This criterion can be straightforwardly generalized to include even-parity states. The transformed Wigner function (7) of the transmitted CS state, regardless of its parity, attains negative values if and only if the channel satisfies (10). The necessary and sufficient condition
$\exists(X,P)\in\mathbb{R}^{2}:W^{\prime}_{\xi,\pm}(X,P)<0\iff f_{X}^{2}f_{P}^{2}-\sigma_{X}\sigma_{P}>0$ (11)
depends only on the parameters of the channel. If not satisfied, negative areas in the transmitted CS state, regardless of its parity or magnitude, can not be preserved by the discussed Gaussian squeezing operations. The conditon (11) is proven within the supplementary material.
@@
Minimizing decoherence across multiple noisy channels
@@
$f_{X}$ $=\sqrt{\eta\eta^{\prime}}\mathrm{e}[-(\gamma+\gamma^{\prime})]\;\text{,}$ (13)
$\sigma_{X}$ $=2(1-\eta^{\prime})\mathrm{e}[-2\gamma^{\prime}_{t}]V^{\prime}+2\eta^{\prime}(1-\eta)\mathrm{e}[-2(\gamma^{\prime}+\gamma_{t})]V\;\text{,}$
$f_{P}$ $=\sqrt{\eta\eta^{\prime}}\mathrm{e}[+(\gamma+\gamma^{\prime})]\;\text{,}$
$\sigma_{P}$ $=2(1-\eta^{\prime})\mathrm{e}[+2\gamma^{\prime}_{t}]V^{\prime}+2\eta^{\prime}(1-\eta)\mathrm{e}[+2(\gamma^{\prime}+\gamma_{t})]V^{\prime}\;\text{,}$
@@
In figure 5 we explore the central negativity of the odd-parity CS state with ${\sqrt{2}\xi=3}$ transmitted through a composite channel (13). Without loss of generality we set the transmittances in both channels identical to keep the visualisation two dimensional. The plots show the best attainable negativity and the optimal pre-squeezing and mid-squeezing rates. We consider a number of asymmetric thermal environments. In figure 5 (a) the dashed lines represent the attainable CN without any protective squeezing operations applied. The solid black line represents the best attainable CN when optimal pre-squeezing and mid-squeezing operations are applied to the CS state. The optimal CN depends only on the symmetric variance of the noise, rather than its asymmetry. The solid black line can be equally obtained for the same channel where both environments are symmetric. The solid black line in figure 5 (b) represents the optimal pre-squeezing rate $\gamma$ obtained when there is no asymmetry in either thermal environment, whereas the colored lines correspond to different asymmetries. We recognize that these rates are constantly shifted from the symmetric case, the offset equal to $\gamma_{t}$ of the first channel. The colored bullets, obtained by adding the respective $\gamma_{t}$ rates to the black line, are used to highlight this observation. Furthermore, in the third plot, figure 5 (c), we observe that the optimal mid-squeezing rates $\gamma^{\prime}$ match the differences ${\gamma^{\prime}\equiv\gamma^{\prime}_{t}-\gamma_{t}}$ between the asymmetries of both environments.
@@
Hilbert-Schmidt distance as a measure of state adaptation
@@ Acknowledgements
The authors acknowledge early discussions with H. Le Jeannic. J.P. acknowledges fruitful discussions with O. Solodovnikova and J. Fiurášek. J.P. acknowledges using the computational cluster at the Department of Optics at Palacký University and using several open-source software libraries harris2020,virtanen2020,dalcin2021,hunter2007,meurer2017 in the computation and subsequent evaluation of presented results.
@@
Department of Optics, Palacký University, 17. listopadu 1192/12, 771 46 Olomouc, Czech Republic
Petr Marek [Uncaptioned image]
Department of Optics, Palacký University, 17. listopadu 1192/12, 771 46 Olomouc, Czech Republic
Julien Laurat [Uncaptioned image]
Laboratoire Kastler Brossel, Sorbonne Université, CNRS, ENS-Université PSL, Collège de France,
4 place Jussieu, 75005 Paris, France
Radim Filip [Uncaptioned image]
Department of Optics, Palacký University, 17. listopadu 1192/12, 771 46 Olomouc, Czech Republic
Conditions for presence of negativity in Wigner functions
It convenient to rewrite the complex-exponential form of the output Wigner function of the CS state into its manifestly real form,
@@
where the factors $m_{\xi,\pm}$ stand for the normalization, given as
$m_{\xi,\pm}^{-1}=2\pi\sqrt{V_{X}V_{P}}\left(1\pm\exp(-X_{0}^{2}-P_{0}^{2})\right)\;\text{.}$ (S.2)
The coefficients $f_{X},f_{P},\sigma_{X},\sigma_{P},V_{X}$ and $V_{P}$ are all real and positive.
Let us begin with the sufficient and necessary condition for the existence of any negative area in the Wigner function of the odd-parity CS state. It is only natural to consider the central negativity in this case. The formula (S.1) simplifies into
@@
Direct comparison of the arguments within the exponential functions reveals the condition
$-X_{0}^{2}\left(\frac{f_{X}f_{P}-\sigma_{X}\sigma_{P}}{V_{X}V_{P}}\right)-P_{0}^{2}\left(\frac{f_{X}f_{P}-\sigma_{X}\sigma_{P}}{V_{X}V_{P}}\right)<0\iff f_{X}^{2}f_{P}^{2}-\sigma_{X}\sigma_{P}>0\;\text{.}$ (S.5)
The central point of the odd-parity CS state will be negative if and only if ${f_{X}^{2}f_{P}^{2}-\sigma_{X}\sigma_{P}>0}$. Notably this inequality does not depend on the state. It depends solely on the channel. This concludes the proof for the odd-parity state. $\blacksquare$
The derivation of the condition for even-parity states is slightly more involved. We are looking for the existence of negative regions. This allows us to be somewhat liberal in the manipulation of the expression (S.1).
@@
±$\pm\cos\left(2\left[\frac{f_{P}}{V_{P}}X_{0}P^{\prime}-\frac{f_{X}}{V_{X}}P_{0}X^{\prime}\right]\right)\equiv-\left\lvert\cos\left(2\left[\frac{f_{P}}{V_{P}}X_{0}P^{\prime}-\frac{f_{X}}{V_{X}}P_{0}X^{\prime}\right]\right)\right\rvert\;\text{,}$ (S.6)
and, similarly, infinitely many $(X^{\prime},P^{\prime})$ pairs that also satisfy
$\cosh\left(2\left[\frac{f_{X}}{V_{X}}X_{0}X^{\prime}+\frac{f_{P}}{V_{P}}P_{0}P^{\prime}\right]\right)\equiv 1\;\text{.}$ (S.7)
The points ${(X^{\prime},P^{\prime})}$ satisfying both relations simultaneously can be determined analytically as
$X^{\prime}_{k}$ $=-\frac{V_{X}}{f_{X}}\frac{P_{0}}{X_{0}^{2}+P_{0}^{2}}\frac{2k+1}{2}\pi\;\text{,}$ (S.8)
$P^{\prime}_{k}$ $=+\frac{V_{P}}{f_{P}}\frac{X_{0}}{X_{0}^{2}+P_{0}^{2}}\frac{2k+1}{2}\pi\;\text{,}$
by solving a pair of parametric linear equations with respect to $X^{\prime}$ and $P^{\prime}$,
$\frac{f_{X}}{V_{X}}X_{0}X^{\prime}+\frac{f_{P}}{V_{P}}P_{0}P^{\prime}$ $=0\;\text{,}$ (S.9)
$\frac{f_{P}}{V_{P}}X_{0}P^{\prime}-\frac{f_{X}}{V_{X}}P_{0}X^{\prime}$ $=\frac{2k+1}{2}\pi\;\text{.}$
where the parameter ${\lvert k\rvert\in\mathbb{N}}$ embodies the periodicity of the cosine function. The Wigner function (S.1) simplifies into (S.3) for these points and the proof reduces into the odd-parity case. $\blacksquare$
Closed form formula for the Hilbert-Schmidt distance
@@
$Q_{\xi}$ $=2\pi\iint\limits_{\mathbb{R}^{2}}W_{\xi,+}(X,P)W_{\xi,-}(X,P)\mathrm{d}{X}\mathrm{d}{P}$